OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x3 + 5x2 - 4x - 20 A. x = -5, x = 4 B. x = -2, x = 2 C. x = -5, x = -2, x = 2 D. x = 5, x = -2, x = 2

OpenStudy (anonymous):

Factor by grouping: \[x^3+5x^2-4x-20=0\\ x^2(x+5)-4(x+5)=0\\ (x^2-4)(x+5)=0\]

OpenStudy (anonymous):

ok cool so whats next

OpenStudy (anonymous):

Do you know how to solve this? If you keep factoring, you get \[(x+2)(x-2)(x+5)=0\] From here, you split up into three equations and solve for x: \[\begin{matrix}x+2=0,&x-2=0,&x+5=0\end{matrix}\]

OpenStudy (anonymous):

is it C @SithsAndGiggles

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

wow thanks i think that deserves a medal lol @SithsAndGiggles

OpenStudy (anonymous):

You're welcome